### v3

@@ 4 Application of the POf to composite problems
In this context, adding a constraint $f(y)=x$, with any $x\in\mathbb{X}$ fixed, simplifies Problem ($\mathbf{P}_{\texttt{i}}$). Indeed, $f(y)=x$ is a smooth equality constraint and $\sigma(f(y))=\sigma(x)$ and $\varepsilon(f(y))=\varepsilon(x)$ are constant. Accordingly, the POf partitions $\mathbb{Y}=\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$ with $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$ for all $x\in\mathbb{X}$. Then, either Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) is infeasible or, by assumption, a global solution $\gamma(x)$ may be computed. Thus, Reformulation ($\mathbf{P}_{\texttt{r}}$) is a low-dimensional blackbox problem where, given any $x\in\mathbb{X}$, we try to compute $\gamma(x)$ and $\sigma(x)$ and $\varepsilon(x)$ on-the-fly, and we return either $\Phi(x)\triangleq\widetilde{\varphi}(\gamma(x),\sigma(x))+\varepsilon(x)$ if $\gamma(x)$ exists or $\Phi(x)\triangleq+\infty$ otherwise. We then use the DFPOm to solve Reformulation ($\mathbf{P}_{\texttt{r}}$) and deduce a (possibly, generalized) local solution to Problem ($\mathbf{P}_{\texttt{i}}$). Also, remark that this partition yields $\chi\triangleq f$ so the continuity of $\chi$ naturally holds.
This section solves four examples using the DFPOm. In Section 4.1, we solve the example from Section 1.1. In Section 4.2, $f$ returns the radius of the polar coordinates and $\gamma$ is easily tractable. In Section 4.3, $f$ returns a nonlinear combination of all variables and the expression of $\gamma$ may be calculated. In Section 4.4, $f$ returns two nonlinear combinations of all variables, and $\gamma$ cannot be expressed in closed form but Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) is solvable with a dedicated algorithmic method. In all cases, we implement the cDSM as described in Remark 5. These examples show the variety of what Theorem 2 may guarantee. In Sections 4.1 and 4.3, Theorem 2 ensures the optimality of $\gamma(x^{*})$ only in the generalized local sense, although $\gamma(x^{*})$ is a usual local solution in Section 4.1. In Sections 4.2 and 4.4, Theorem 2 successfully claims the usual local optimality of $\gamma(x^{*})$.
@@ 4.1 Mono-variable noise
pictured in Figure 2 (right) and Figure 3 (left). It follows that Reformulation ($\mathbf{P}_{\texttt{r}}$) minimizes
@@ 4.4 Bi-dimensional noise and oracle provided by an algorithmic method
so $M(x)$ is obtainable via a dichotomic search and the singleton $\{T(x)\}$ follows. In practice, running sufficiently many iterations of the dichotomic search provides an accurate approximation of $M(x)$. We define $M^{0}_{\mathrm{inf}}\triangleq\left\lfloor{M(x)}\right\rfloor$ and $M^{0}_{\mathrm{sup}}\triangleq M^{0}_{\mathrm{inf}}+1$, and $\ell\triangleq 0$, and we iterate the dichotomic search on $M^{\ell}_{\mathrm{inf}},M^{\ell}_{\mathrm{sup}}]$ while $M^{\ell}_{\mathrm{sup}}-M^{\ell}_{\mathrm{inf}}>2^{-30}$. This allows to define $M(x)\approx\hat{M}(x)\triangleq\frac{1}{2}(M^{\ell}_{\mathrm{inf}}+M^{\ell}_{\mathrm{sup}})$, and then $T(x)\approx\hat{T}(x)\triangleq\frac{1}{2}(\min I_{x}(\hat{M}(x))+\max I_{x}(\hat{M}(x)))$, and finally $\gamma(x)\approx\hat{\gamma}(x)\triangleq s_{x}(\hat{T}(x))$.
@@
Figure 9: Partition sets of $\mathbb{Y}(x)$ and locus of $\gamma$ in the context of Section $4.$
Figure 10: Functions (left) $\varepsilon$, (centre) $M$ and (right) $\Phi$ in Section $4.$ In this section $\Phi=\varepsilon+M$.
@@
Table 4: Results for eight instances of Algorithm 2 in the context of Section $4.$ In this table, $\hat{x}^{k}\triangleq\lVert{x^{k}}\rVert_{\infty}$.
@@ 5 Gain of performance provided by the DFPOm
In this section, we tackle composite greybox problems where the POf allows for Reformulation ($\mathbf{P}_{\texttt{r}}$) with $\dim(\mathbb{X})$ small (equal to the number of inputs of the blackboxes $\sigma$ and $\varepsilon$) even though $\dim(\mathbb{Y})$ is possibly large. We show that the DFPOm is efficient in this context. We alter the problems from Section 4 so that $\dim(\mathbb{Y})\approx 100$ while $\dim(\mathbb{X})\leq 10$. Let us stress that our choice to consider $\dim(\mathbb{Y})\approx 100$ at most stands only for limited computational effort: $\dim(\mathbb{Y})\approx 100$ is already larger than what many DFO algorithms tackle, and our examples could be adapted to match any $0<\dim(\mathbb{X})\leq\dim(\mathbb{Y})\leq+\infty$ (we even have $\dim(\mathbb{X})\leq 4\ll\dim(\mathbb{Y})=+\infty$ in [11]).
@@
In most examples below, computing $\varphi(y)$, $\gamma(x)$ and $\Phi(x)$ is very fast for all $y\in\mathbb{Y}$ and $x\in\mathbb{X}$. This likely differs from real cases, where either $\varphi(y)$ may be computed by an expensive computer program and $\gamma(x)$ may be obtained by numerically solving an optimization subproblem. Thus, to establish a realistic comparison between the solvers, we proceed as follows. For each problem, we consider that the cost to compute $\varphi(y)$ is $1$ unit, identical for all $y\in\mathbb{Y}$, and that the computation of $\Phi(x)$ costs $1+\tau$ units, identical for all $x\in\mathbb{X}$ and where $\tau\geq 0$ is fixed. The value $\tau$ captures the relative cost to evaluate $\gamma$ versus those to evaluate $\varphi$, so the computational cost of $\Phi(x)$ models the additional cost required to first compute $\gamma(x)$ before computing $\varphi(\gamma(x))$. Then, for each method, we track the sequence of all evaluated points. This consists in evaluations of $\varphi$ for solvers tackling Problem ($\mathbf{P}_{\texttt{i}}$) directly, and of $\Phi$ for our method focusing on Reformulation ($\mathbf{P}_{\texttt{r}}$) instead. Finally, for each method we plot the graph representing the lowest objective value found within a budget of $c$ units, for all $c\in\mathbb{R}_{+}$.
@@
We compare the DFPOm against the DFO solvers NOMAD [7] and PRIMA [35] solving Problem ($\mathbf{P}_{\texttt{i}}$) directly. We observe that the DFPOm significantly outperforms both NOMAD and PRIMA, unless $\tau$ is prohibitively large. Note that this does not discredit these solvers: as we discuss in Section 2, they are not designed for large-dimensional contexts, and others DFO techniques [19, 21, 22, 23, 26] may also perform poorly in contexts suited for the POf. Similarly, our results do shred light on our instance of the cDSM. What makes the DFPOm efficient is the sheer dimension reduction provided by the POf.
@@
We follow the extreme barrier approach [4] that minimizes $\overline{\varphi_{\Omega}}$ over $\mathbb{Y}$, with relevant box constraints. We run NOMAD with default parameters, PRIMA with the BObyQA algorithm, and the cDSM as in Remark 5. We give to NOMAD and PRIMA six random starting points $(y^{0}_{\ell})_{\ell\in\llbracket 1,6\rrbracket}$, and we give to our cDSM the starting points $(\chi(y^{0}_{\ell}))_{\ell\in\llbracket 1,6\rrbracket}$. In this section, we denote by $\mathds{1}\triangleq(1,\dots,1)$.
@@
The DFPOm outperforms NOMAD and PRIMA tackling Problem ($\mathbf{P}_{\texttt{i}}$) directly. Our cDSM closely approaches the global minimizer while NOMAD and PRIMA remain in an exploratory phase that returns poor results, and the computational budget is comparable when $\tau=10$.
@@ 5.3 Noise affected by a nonlinear combination of all 100 variables
Similarly to Sections 5.1 and 5.2, Reformulation ($\mathbf{P}_{\texttt{r}}$) is a DFO problem fixing the value of a single combination of all the variables, and the POf fixes the $100$ variables of the original problem accordingly. Even with $\tau=100$, the POf provides a gain of performance. NOMAD and PRIMA solving the original problem both fail to significantly improve their initial incumbent solution and, in the second case, both solvers behave as if their initial incumbent solution cannot be improved. Within a comparable budget, the cDSM in our DFPOm converges towards the global minimizer on all but one case. Yet, in this case, first we observe that this solution remains $1000$ times better than those returned by NOMAD and PRIMA, and second we claim that the poor performance results from our naive implementation: even if Reformulation ($\mathbf{P}_{\texttt{r}}$) is $1$-dimensional, it is preferable to use a globalization strategy because the interval of possibly relevant values is large. Our instance lacks one since it lacks a search step.
@@ 5.4 Ten-dimensional noise with non-analytical oracle and 100 variables
Figure 14 shows a comparison between the different strategies, where for tractability we actually minimize $\hat{\Phi}$ instead of $\Phi$. Even if the problem we actually solve slightly differs from the true problem, we observe results similar to those in Sections 5.1, 5.2 and 5.3. Reformulation ($\mathbf{P}_{\texttt{r}}$) is a $10$-dimensional DFO problem fixing the value of the important combinations of all the variables, and the POf highlights how to fix the $100$ variables of Problem ($\mathbf{P}_{\texttt{i}}$) accordingly. Using our cDSM, the DFPOm returns solutions $1E2$ times better than those returned by NOMAD and $1E7$ times better than those returned by PRIMA, and the computational budget is similar if we allow $\tau\approx 10$ at most. The value $\tau=10$ is low, but we could consider a higher value by using a better algorithm. For example, a test with NOMAD solving the reformulated problem (instead of our naive instance of cDSM) provides similar graphs with $\tau=30$, represented on Figure 15.
@@ 6.1 Comments on the POf and the DFPOm
The POf is a reformulation framework that applies to a large class of problems, including nonconvex, discontinuous and unbounded problems. The idea is to pinpoint the difficult aspect in Problem ($\mathbf{P}_{\texttt{i}}$), and to partition the variables space so that this difficulty is transparent in each partition set. Then, the POf defines the oracle function $\gamma$ (assumed computable) that is assumed to return a solution to the problem restricted to any partition set (which is easier to solve than the unrestricted problem). Thus, the POf is not a DFO framework by itself, but rather a way to reformulate a problem with localized difficulties so that Reformulation ($\mathbf{P}_{\texttt{r}}$) tackles only these difficulties. Then, Reformulation ($\mathbf{P}_{\texttt{r}}$) is manageable using DFO algorithms, as we formalize as the DFPOm. This method is theoretically grounded (as formalized by Theorem 2), since Theorem 1 ensures that solutions to Reformulation ($\mathbf{P}_{\texttt{r}}$) lead to (possibly, generalized) local solutions to Problem ($\mathbf{P}_{\texttt{i}}$), and Proposition 1 ensures that Algorithm 1 indeed solves Reformulation ($\mathbf{P}_{\texttt{r}}$).
@@
A stringent practical aspect of the POf is the oracle function $\gamma$. For all $x\in X$, Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) must have a computable global solution. Ensuring this requirement is challenging and likely problem-dependent. In some cases, $\gamma$ is analytically tractable. However, the POf does not need that $\gamma$ has a closed-form expression. In some of our examples, $\gamma$ is defined pointwise as, for all $x\in X$, the output of some numerical method solving Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) with high confidence to identify a global solution (e.g., by solving a simple smooth subproblem). Our experiments in our former work [11] even show that the DFPOm remains effective when $\gamma(x)$ consists in the output of some solver tackling Subproblem ($\mathbf{P}_{\texttt{s}}(x)$), even if the solver possesses only guarantees to identify a point satisfying necessary optimality conditions. These numerical studies indicate that such approximations are sufficient in practice to reap the benefits of the POf, while our theory justifies the methodology when $\gamma$ returns global solutions. This remains restrictive, but some applications already fit in this framework (as discussed in Sections 1.2, 4 and 5), and we are working on possible ways to significantly relax this requirement (see Section 6.2).
Our DFPOm solves Reformulation ($\mathbf{P}_{\texttt{r}}$) using a DFO algorithm with a covering step. We consider the cDSM (Algorithm 2) in our numerical experiments, but any DFO algorithm may be enhanced with a covering step, as discussed in Section 3.2 and [3]. The choice of the most appropriate DFO algorithm is likely to depend on $\Phi$. In general, Reformulation ($\mathbf{P}_{\texttt{r}}$) has little specificity that differentiates it from a generic blackbox optimization problem, so there is no reason a priori to favour a DSM instead of another class of methods. In particular, if $\dim(\mathbb{X})$ is too large, a DSM may not be suited to solve Reformulation ($\mathbf{P}_{\texttt{r}}$) while another DFO may be more appropriate. However, to our best knowledge, no popular DFO solver implements some algorithms with a covering step yet. Then, in practical settings, all fast solvers we may consider are only guaranteed to return a point satisfying necessary optimality conditions (instead of local optimality). Thus, Theorem 2 is not yet compatible with most DFO solvers. Nevertheless, [3] remarks that the covering step is likely optional in practice. We second this observation in Section 5.4, as solving Reformulation ($\mathbf{P}_{\texttt{r}}$) using NOMAD instead of the cDSM provides better results, even though NOMAD does not implement a covering step.
The usual case of application of the DFPOm is $\dim(\mathbb{Y})\geq 2$ and $0<\dim(\mathbb{X})\ll\dim(\mathbb{Y})$. It is advised to ensure that $\dim(\mathbb{X})\leq 50$ to avoid bad performance when solving Reformulation ($\mathbf{P}_{\texttt{r}}$) with the cDSM [6, Section 1.4]. Although our examples in Section 5 illustrate cases with $\dim(\mathbb{Y})=101$ at most, we stress that the performance of the POf depends only on $\dim(\mathbb{X})$ and on the computation time related to $\gamma$. Hence, $\dim(\mathbb{Y})$ may be possibly infinite, as in [11]. Moreover, the smaller $\dim(\mathbb{X})$ is and the more efficient the DFO solver solving Reformulation ($\mathbf{P}_{\texttt{r}}$) is, the larger the constant $\tau$ from Section 5 (the relative cost to evaluate $\gamma$ versus those to evaluate $\varphi$) can be. As a rule of thumb, if we can estimate the number $\mathbf{N}^{\texttt{ref}}$ of points in $\mathbb{X}$ required by the DFO solver to solve Reformulation ($\mathbf{P}_{\texttt{r}}$), and $\mathbf{N}^{\texttt{ini}}$ of points in $\mathbb{Y}$ required by a dedicated solver to solve Problem ($\mathbf{P}_{\texttt{i}}$) directly, then the DFPOm is useful when $\mathbf{N}^{\texttt{ref}}\tau<\mathbf{N}^{\texttt{ini}}$. We also remark that the way the partition is indexed has a direct impact on the behaviour of $\Phi$. The solving process may be substantially simplified by a re-parametrization of $\mathbb{X}$ that avoid concentrations of discontinuities or mitigate local variations of $\Phi$.

### v5

@@ 5 Application of the POf to composite problems
In this context, adding a constraint $f(y)=x$, with any $x\in\mathbb{X}$ fixed, simplifies Problem ($\mathbf{P}_{\texttt{ini}}$). Indeed, $f(y)=x$ is a smooth equality constraint and $\sigma(f(y))=\sigma(x)$ and $\varepsilon(f(y))=\varepsilon(x)$ are constant. Accordingly, the POf partitions $\mathbb{Y}=\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$ with $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$ for all $x\in\mathbb{X}$ (note that this partition yields $\chi\triangleq f$ so the continuity of $\chi$ naturally holds). Then, either Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) is infeasible or, by assumption, a global solution $\gamma(x)$ may be computed. Thus, Reformulation ($\mathbf{P}_{\texttt{ref}}$) is a low-dimensional blackbox problem where, given any $x\in\mathbb{X}$, we try to compute $\gamma(x)$ and $\sigma(x)$ and $\varepsilon(x)$ on-the-fly, and we return either $\Phi(x)\triangleq\widetilde{\varphi}(\gamma(x),\sigma(x))+\varepsilon(x)$ if $\gamma(x)$ exists or $\Phi(x)\triangleq+\infty$ otherwise. We then use the DFPOm to solve Reformulation ($\mathbf{P}_{\texttt{ref}}$) and deduce a (possibly, generalized) local solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
This section solves four examples using the DFPOm. In Section 5.1, we solve the example from Section 1.1. In Section 5.2, $f$ returns the radius of the polar coordinates and $\gamma$ is easily tractable. In Section 5.3, $f$ returns a nonlinear combination of all variables and the expression of $\gamma$ may be calculated. In Section 5.4, $f$ returns two nonlinear combinations of all variables, and $\gamma$ cannot be expressed in closed form but Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) is solvable with a dedicated algorithmic method. These examples show the variety of what Theorem 2 may guarantee. In Sections 5.1 and 5.3, Theorem 2 ensures the optimality of $\gamma(x^{*})$ only in the generalized local sense, although $\gamma(x^{*})$ is a usual local solution in Section 5.1. In Sections 5.2 and 5.4, Theorem 2 successfully claims the usual local optimality of $\gamma(x^{*})$. In all cases, we implement the DFPOm with the cDSM as described in Remark 5, so we solve Reformulation ($\mathbf{P}_{\texttt{ref}}$) numerically up to some stopping criterion and the DFPOm outputs an approximation of a solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
@@ 5.1 Mono-variable noise
pictured in Figure 2 (right) and Figure 3 (left). It follows that Reformulation ($\mathbf{P}_{\texttt{ref}}$) minimizes
@@ 5.4 Bi-dimensional noise and oracle provided by an algorithmic method
so $M(x)$ is obtainable by dichotomic search and the singleton $\{T(x)\}$ follows. In practice, running sufficiently many iterations of the dichotomic search provides an accurate approximation of $M(x)$. We define $M^{0}_{\mathrm{inf}}\triangleq\left\lfloor{M(x)}\right\rfloor$ and $M^{0}_{\mathrm{sup}}\triangleq M^{0}_{\mathrm{inf}}+1$, and $\ell\triangleq 0$, and we iterate the dichotomic search on $M^{\ell}_{\mathrm{inf}},M^{\ell}_{\mathrm{sup}}]$ while $M^{\ell}_{\mathrm{sup}}-M^{\ell}_{\mathrm{inf}}>2^{-30}$. This allows to define $\hat{M}(x)\triangleq\frac{1}{2}(M^{\ell}_{\mathrm{inf}}+M^{\ell}_{\mathrm{sup}})\approx M(x)$, and then $\hat{T}(x)\triangleq\frac{1}{2}(\min I_{x}(\hat{M}(x))+\max I_{x}(\hat{M}(x)))\approx T(x)$, and finally $\hat{\gamma}(x)\triangleq s_{x}(\hat{T}(x))\approx\gamma(x)$.
@@
Table 4: Results for eight instances of Algorithm 2 in the context of Section 5.4. In this table, $\hat{x}^{k}\triangleq\lVert{x^{k}}\rVert_{\infty}$.
Figure 9: Partition sets of $\mathbb{Y}(x)$ and locus of $\gamma$ in the context of Section 5.4.
Figure 10: Functions (left) $\varepsilon$, (centre) $M$ and (right) $\Phi$ in Section 5.4. In this section $\Phi=\varepsilon+M$.
@@ 6 Gain of performance provided by the DFPOm
In this section, we tackle composite greybox problems where the POf allows for Reformulation ($\mathbf{P}_{\texttt{ref}}$) with $\dim(\mathbb{X})$ small (equal to the number of inputs of the blackboxes $\sigma$ and $\varepsilon$) even though $\dim(\mathbb{Y})$ is possibly large. We show that the DFPOm is efficient in this context. We alter the problems from Section 5 so that $\dim(\mathbb{Y})\approx 100$ while $\dim(\mathbb{X})\leq 10$. Let us stress that our choice to consider $\dim(\mathbb{Y})\approx 100$ at most stands only for limited computational effort: $\dim(\mathbb{Y})\approx 100$ is already larger than what many DFO algorithms tackle, and our examples could be adapted to match any $0<\dim(\mathbb{X})\leq\dim(\mathbb{Y})\leq+\infty$ (we even have $\dim(\mathbb{X})\leq 4\ll\dim(\mathbb{Y})=+\infty$ in [10]).
@@
In our examples below, computing $\varphi(y)$, $\gamma(x)$ and $\Phi(x)$ is very fast for all $y\in\mathbb{Y}$ and $x\in\mathbb{X}$. This likely differs from real cases, where $\varphi(y)$ may be computed by an expensive computer program and $\gamma(x)$ may be obtained by numerically solving an optimization subproblem. Thus, to establish a realistic comparison between the solvers, we proceed as follows. For each problem, we consider that the cost to compute $\varphi(y)$ is $1$ unit, identical for all $y\in\mathbb{Y}$, and that the computation of $\Phi(x)$ costs $1+\tau$ units, identical for all $x\in\mathbb{X}$ and where $\tau\geq 0$ is fixed. The value $\tau$ captures the relative cost to evaluate $\gamma$ versus those to evaluate $\varphi$, so the computational cost of $\Phi(x)$ models the additional cost required to first compute $\gamma(x)$ before computing $\varphi(\gamma(x))$. Then, for each method, we track the sequence of all evaluated points. This consists in evaluations of $\varphi$ for solvers tackling Problem ($\mathbf{P}_{\texttt{ini}}$) directly, and of $\Phi$ for our method focusing on Reformulation ($\mathbf{P}_{\texttt{ref}}$) instead. Finally, for each method we plot the graph representing the lowest objective value found within a budget of $c$ units, for all $c\in\mathbb{R}_{+}$.
@@
We compare the DFPOm against the DFO solvers Nomad [7] and Prima [39] solving Problem ($\mathbf{P}_{\texttt{ini}}$) directly. We observe that the DFPOm significantly outperforms both Nomad and Prima, unless $\tau$ is prohibitively large. Note that this does not discredit these solvers: as we discuss in Section 2, they are not designed for contexts suited for the POf, and others DFO techniques [20, 24, 22, 23, 29] may also perform poorly in these contexts. Similarly, our results do not shred light on our instance of the cDSM in the DFPOm. What makes the DFPOm efficient is the sheer dimension reduction provided by the POf.
@@
We follow the extreme barrier approach [4] that minimizes $\overline{\varphi_{\Omega}}$ over $\mathbb{Y}$, with relevant box constraints. We run Nomad with default parameters, Prima with the BObyQA algorithm, and the DFPOm as in Remark 5. We give to Nomad and Prima six random starting points $(y^{0}_{\ell})_{\ell\in\llbracket 1,6\rrbracket}$, and we give to our cDSM the starting points $(\chi(y^{0}_{\ell}))_{\ell\in\llbracket 1,6\rrbracket}$. In this section, we denote by $\mathds{1}\triangleq(1,\dots,1)$.
@@
The DFPOm outperforms Nomad and Prima tackling Problem ($\mathbf{P}_{\texttt{ini}}$) directly. Our cDSM closely approaches the global minimizer while Nomad and Prima remain in an exploratory phase that returns poor results, and the computational budget is comparable when $\tau=10$.
@@ 6.3 Noise affected by a nonlinear combination of all 100 variables
Similarly to Sections 6.1 and 6.2, Reformulation ($\mathbf{P}_{\texttt{ref}}$) is a DFO problem fixing the value of a single combination of all the variables, and the POf fixes the $100$ variables of the original problem accordingly. Even with $\tau=100$, the POf provides a gain of performance. Nomad and Prima solving the original problem both fail to significantly improve their initial incumbent solution and, in the second case, both solvers behave as if their initial incumbent solution cannot be improved. Within a comparable budget, the cDSM in our DFPOm converges towards the global minimizer on all but one case. Yet, in this case, first we observe that this solution remains $1000$ times better than those returned by Nomad and Prima, and second we claim that the poor performance results from our naive implementation: even if Reformulation ($\mathbf{P}_{\texttt{ref}}$) is $1$-dimensional, it is preferable to use a globalization strategy because the interval of possibly relevant values is large. Our instance lacks one since it lacks a search step.
@@ 6.4 Ten-dimensional noise with non-analytical oracle and 100 variables
Figure 14 shows a comparison between the different strategies, where for tractability we actually minimize $\hat{\Phi}$ instead of $\Phi$. Even if the problem we actually solve slightly differs from the true problem, we observe results similar to those in Sections 6.1, 6.2 and 6.3. Reformulation ($\mathbf{P}_{\texttt{ref}}$) is a $10$-dimensional DFO problem fixing the value of the important combinations of all the variables, and the POf highlights how to fix the $100$ variables of Problem ($\mathbf{P}_{\texttt{ini}}$) accordingly. Using our cDSM, the DFPOm returns solutions $1E2$ times better than those returned by Nomad and $1E7$ times better than those returned by Prima, and the computational budget is similar if we allow $\tau\approx 10$ at most. The value $\tau=10$ is low, but we could consider a higher value by using a better algorithm. For example, a test with Nomad solving the reformulated problem (instead of our naive instance of cDSM) provides similar graphs with $\tau=30$, represented on Figure 15.
@@ 7.1 Comments on the POf and the DFPOm
The POf is a reformulation framework that applies to several classes of problems, including infinite-dimensional, nonconvex, discontinuous and unbounded problems. The idea is to pinpoint the difficult aspect in Problem ($\mathbf{P}_{\texttt{ini}}$), and to partition the variables space so that this difficulty is transparent in each partition set. Then, the POf defines the oracle function $\gamma$ (assumed computable) that maps any partition set index to a solution to the (simpler) problem restricted to the partition set. Thus, the POf is not a DFO framework by itself, but rather a way to reformulate a problem with localized difficulties so that Reformulation ($\mathbf{P}_{\texttt{ref}}$) tackles only these difficulties. Then, Reformulation ($\mathbf{P}_{\texttt{ref}}$) is manageable using DFO algorithms, as we formalize in the DFPOm. This method is theoretically grounded, as formalized by Theorem 2, since Theorem 1 ensures that solutions to Reformulation ($\mathbf{P}_{\texttt{ref}}$) lead to solutions to Problem ($\mathbf{P}_{\texttt{ini}}$), and Proposition 1 ensures that Algorithm 1 indeed solves Reformulation ($\mathbf{P}_{\texttt{ref}}$).
@@
A stringent practical aspect of the POf is the oracle function $\gamma$. For all $x\in X$, Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) must have a computable global solution. Ensuring this requirement is challenging and likely problem-dependent, but some applications already fit in this framework (as discussed in Sections 1.2, 4, 5 and 6), and we are working on possible ways to significantly relax this requirement (see Section 7.2). In some cases, $\gamma$ is analytically tractable, but the POf does not need that $\gamma$ has a closed-form expression.